OpenStudy (anonymous):

Find the solution set. Separate the two values with a comma: 7x^2+x=6

OpenStudy (anonymous):

7x^2 + x -6 =0 7x^2 +7x -6x -6 =0 7x(x + 1) - 6 (x+1) =0 (7x-6)(x+1)=0 x=6/7,-1